# What If Moderation Didn’t Mean Suppression? A Case for Personalized Content Transformation

### v1

@@ 4.3. Server Architecture and Processing Pipeline
The computational demands of this server processing pipeline are substantial. For example, processing a single Reddit batch can require 50 to 250 sequential external LVLM API calls that must complete within seconds to maintain usability. We achieved real-time performance through asynchronous processing with multi-batch responses, content-based caching, and predictive prefetching. We have described these optimizations and additional design decisions in Appendix˜A.
@@ 4.4. The Intervention Palette
The design of our system is based on a shift from content suppression to transformation. In taking this approach, we faced a design challenge: every intervention has to navigate a trade-off between three competing goals. The first is semantic fidelity, or how faithful the modified content is to the original’s intended meaning. The second is trigger fidelity, or how close the modified distressing element is to its original form. The final goal is perceptual smoothness, or how natural the result of the intervention looks to the user. We designed a palette of transformations, each reflecting a different approach to these trade-offs.
@@
In contrast, semantic modifications prioritize perceptual smoothness. Inpainting(Yu et al., 2018; Li et al., 2022) as an example case, can seamlessly remove an object and reconstruct the background. This is a greater sacrifice of semantic fidelity than occlusion because it actively creates a new, plausible reality where the object never existed, rather than simply occluding part of the original. Another semantic intervention, visual euphemism, explores a different trade-off. It replaces a triggering object with a benign alternative, for instance, substituting spider with a leaf. While this also sacrifices semantic fidelity, the goal is to produce a visually coherent image that we hypothesize is less distressing to the user. To personalize this, users can add additional metadata to their filters or give examples to guide the visual euphemism choices, which may in turn alter the mood of the post.
Third, we also explore how altering an image’s rendering style could offer a middle ground. For example, reducing photorealism with artistic alteration, such as  Pointillism (Clement and Houzé, 1999) can lower the trigger fidelity of a sensitive region while preserving more structural information than occlusion. For these alterations, we borrow the aesthetic language of well-known artistic styles—Pointillism, Cubism, Impressionism, and Studio Ghibli style animation. We intended for each artistic style to not only reduce realism, but also act as an affective vessel that shapes how content is experienced. Impressionism, with its soft brushstrokes and diffused light, often evokes a sense of tranquility, distance, or dream-like calm. Studio Ghibli inspired style, with its rounded and whimsical rendering, lends even serious scenes a tone of cuteness and warmth. Cubism, which fragments forms in abstract planes, introduces a kind of conceptual distance, perhaps inviting interpretation over emotional reaction.
@@
In DIY-MOD, we implemented a representative set of interventions that explore the broader design space of content transformation. Figure˜2 illustrates several examples from our image palette’s three main categories. We list all implemented interventions in Appendix˜D. The range of transformation options presents an important question: which intervention is most appropriate for a specific user and post? Next, we detail the framework we developed to address this question.
@@ 4.5. Intervention Selection Framework
We developed a two-stage pipeline that evaluates interventions through the lens of individual user needs while maintaining real-time performance.
@@ 4.5.1. Evaluation Criteria:
The three dimensions of semantic fidelity, trigger fidelity, and perceptual smoothness characterize the design space, but do not, on their own, determine which intervention is best suited for a given scenario. That choice depends on how a user experiences the transformation applied to a content. To guide this selection and evaluate outcomes, we developed a framework that decomposes intervention quality into four complementary dimensions (Table˜1), assessed through LLM-as-judge (Zheng et al., 2023; Liu et al., 2023):
Table 1. Evaluation Framework. All dimensions assess both original and transformed content. Items within parenthesis indicate additional inputs required for personalized evaluation.
Dimension Assesses
Transformation Seamlessness Is the output perceptually smooth for images and coherent for text?
Semantic Fidelity Are the original facts preserved without adding misinformation, a known risk in generative models (Huang et al., 2025; Gao et al., 2024)?
Predicted Emotional Impact
(+ user context, filter metadata)
Does it address this specific user’s safety needs and sensitivities?
Transformation Appropriateness
(+ user context)
Does the transformation respect the inherent gravity or cultural significance of the content?
The final two dimensions function as our “limited user model”, transforming objective measures into personalized assessments. Predicted Emotional Impact synthesizes the user’s context—filter description, sensitivity level, chat history, and metadata preferences—with the intervention’s effectiveness at reducing trigger exposure. This predicts whether the transformation adequately addresses the user’s specific sensitivities.
Transformation Appropriateness assesses the degree to which a transformation might trivialize or distort themes that individuals or broader culture consider sensitive. For instance, applying Studio Ghibli style to war imagery or racial violence would score poorly on this dimension. The evaluator subtracts any trivialization score from the overall intervention rating, penalizing contextually inappropriate interventions.
@@ 4.5.2. Two-Stage Selection Pipeline
Stage 1: Predictive Pruning. The Pruner performs multi-task inference(Son et al., 2024) that first checks whether content matches any user filters. For matching content, it predicts which interventions would most likely succeed for that specific user. We provide the Pruner LVLM with detailed descriptions of each intervention including trade-offs, advantages, disadvantages, and ideal use cases. We also supply our four scoring dimensions and instruct the Pruner to evaluate how each intervention would score. This approach leverages the Pruner to approximate the full scoring process without generating actual transformations and produces a ranked list from which we select the top-$K$ candidates (typically $K=3$).
Stage 2: Generation and Scoring. DIY-MOD applies the top-$K$ candidates to the content, in parallel. Obfuscation techniques are applied locally using Python Pillow (Clark, 2015) library. Semantic modifications and stylistic alterations use Gemini’s image generation API (Google, 2025). The Scorer evaluates each candidate transformation using the evaluation dimensions (Table˜1). We implement two guardrails in the scoring process. First, the Scorer receives both the original and transformed content, and is instructed to penalize misapplied interventions and failures. Second, we employ two-stage prompting (Wang et al., 2024) that improve judge-VLM’s performance and results in more reliable and consistent evaluations. The Scorer first performs chain-of-thought analysis along our evaluation dimensions given the original content, transformed candidate, and user context. Based on this free-form response, it then produces a final numerical score. These scoring operations run in parallel across all $K$ candidates.
The Scorer personalizes evaluation using three user contexts: the filter description, the self-assigned sensitivity level (1-5), and the complete chat history from filter creation. The scores are summed and the highest-scoring intervention is selected as the winner.
@@ 5.1. Methods
Data Collection and Analysis: We transcribed all think-aloud data and interviews, then analyzed them using reflexive thematic analysis (Braun and Clarke, 2019) to identify patterns in user experiences and system interactions. During consent, we explicitly warned participants they might encounter content matching their stated sensitivities and emphasized they could pause, skip content, or withdraw at any point.
@@ 6.1. Method
As participants scrolled through the dual feed, they encountered two distinct, modified versions of each post with text and image transformation applied (details in  Section˜C.2). For each pair, they were asked to select the version that better met their safety needs and explain their reasoning. The two modifications were chosen systematically: one was the intervention ranked highest by DIY-MOD’s selection framework, while the other was randomly selected from the framework’s other high-ranking alternatives (ranks 2-5). To mitigate positional bias, the on-screen placement (left or right) of these two versions was randomized for each post. While the primary task involved comparing the two modified versions, the original unmodified content was also available for context. Participants could request to view it or have it described verbally if needed. The study lasted approximately 30 minutes for each participant.
@@ 6.2.3. Quantitative Alignment with User Preferences
To complement our qualitative findings, we modeled the probability that participants would select the system’s highest-ranked intervention over a nearby alternative—the randomly selected intervention from ranks two through five. We use an intercept-only model to predict our dependent variable, System Choice Favored, which is a binary variable equal to 1 if the participant selected the system’s choice, and 0 otherwise. We use the glmer function from the lme4 package(Bates et al., 2015) in R. For image interventions, the estimated intercept was positive and statistically significant (log-odds = 0.53, $p=.033$), corresponding to a 63% probability that participants chose the system’s top candidate over the close contenders. For text interventions, the alignment was stronger (log-odds = $1.33$, p<.001{p}<.001), corresponding to a 79.1% probability.
@@ 7.1. Transformation Enables Access, Not Avoidance
Empowering users to take moderation decisions into their own hands might raise concerns about whether this agency can trap them in filter bubbles(Pariser, 2011). DIY-MOD can support a different dynamic. By transforming how content appears rather than removing it entirely, the system enables users to engage with content they would otherwise avoid completely.
@@
At the same time, this approach surfaces deeper tensions: Who decides what users should or should not engage with, and why does that decision need to be made in a top-down, prescriptive manner? The First Amendment is often framed as a protection of speech, but legal scholars have emphasized that it also safeguards the right to freedom of listening and the right against compelled listening (Corbin, 2009). While these protections do not extend to platforms as private entities, the principles that they represent remain relevant. Our studies show that content that appears benign to most can be experienced as harmful by others. This variability complicates any universal judgment about what content is “safe” or “appropriate”—at least outside the categories codified by law. And it raises a critical question: Who gets to decide which experiences of harm are valid and which users are entitled to protection?
@@ 7.2. The Case for Platform-Integrated Personalization
While DIY-MOD demonstrates personalized moderation is viable as middleware (Fukuyama et al., 2021), its architecture requires intercepting, externally processing, and then scoring generated transformations. These steps introduce unavoidable latency and complicate implementation.
Platforms however, could avoid these bottlenecks entirely. With direct content access, computational resources, and existing user behavioral signals, they could implement personalized moderation far more efficiently. Platforms already use this infrastructure for ad targeting and safety interventions like suicide prevention (Card, 2018; Meta, 2022). The same systems could also power consensual, user-directed personalized content transformation.
@@
Our approach gives rise to a natural question: Which kinds of content should be addressed through platform-led moderation, and which might be better suited to user-driven transformation? Platforms are widely expected to bear responsibility for addressing harms that have broad, collective consequences such as coordinated manipulation campaigns or abuse like hate speech or targeted harassment that fosters hostile environments. Yet even here, the boundaries are far from settled. The same activity can be cast as harassment or as activism, as disinformation or as political dissent. Platforms’ judgments about what to enforce are never neutral. They are shaped by political pressures, business incentives, and cultural norms (Díaz and Hecht-Felella, 2021; Jahanbakhsh, 2023).
@@ 7.4. Application to Civic Discourse
This transformation-based approach may also hold promise for other types of content not brought up by our study participants. One can imagine its application to civic or ideological discourse: when valuable perspectives are obscured by hostile or inflammatory framing, repackaging that content (e.g., by softening antagonistic language, reducing confrontational cues, or highlighting shared values), could make it easier to engage with. It is also worth noting that there is no clear consensus that exposure to counter-attitudinal information is always beneficial. In fact, a body of research shows that such exposure, especially when unfiltered or confrontational, can exacerbate polarization and deepen animosity (Törnberg, 2022; Bail, 2022). This calls into question that more exposure is always better, and highlights the need for systems that support constructive engagement, rather than unmediated confrontation (Bail, 2022). Altering the tone or phrasing of civic or political content through transformations like ours may offer a path toward openness without overwhelm.
Prior work on bridging divides has largely focused on identifying and ranking content that resonates across ideological lines (Ovadya and Thorburn, 2022; Stray, 2021). While valuable, such approaches depend on the availability of naturally bridgeable posts. But what if we could make more content bridgeable? This of course, raises important questions about authenticity and message integrity: How much transformation is too much? When does reframing a message become misrepresentation? Future work must address these ethical boundaries.
@@ 7.5. Therapeutic Grounding and Applications of Our Work
DIY-MOD’s design is inspired by exposure therapy: enabling safer engagement with a weakened or modified version of a stimulus, rather than avoidance (Rothbaum and Schwartz, 2002; Abramowitz et al., 2019). This principle is reflected in our palette of interventions. Our transformations are designed to function as weakened versions of the original content, and the ability to easily adjust sensitivity levels and filter durations echoes the concept of graduated exposure.
While DIY-MOD is not a clinical intervention in its current form, future versions could complement therapeutic practices, e.g., by supporting graduated exposure or enabling users to modulate their engagement with distressing content over time, if they so choose.
@@ 7.6.1. Architectural Safeguards:
Using AI to interpret personal sensitivities carries inherent risks, as system failures can expose users to content they explicitly sought to avoid. One way we mitigate these risks is by implementing a two-stage cascade architecture: a pruner and then a scorer VLM-judge evaluate content using a consistent rubric. Such multi-stage evaluation strengthens the reliability of LLM/VLM-judge based pipelines (Badshah and Sajjad, 2024; Gu et al., 2024).
@@
Privacy considerations are fundamental when handling data tied to personal sensitivities. DIY-MOD adopts a “privacy-by-design” approach (Cavoukian et al., 2009) that prioritizes user control and data minimization. It provides full functionality without an account and offers user-controlled portability of filters through an anonymous export/import feature. By default, it uses local-first storage and practices data minimization, sending only public in-feed content for ephemeral analysis. Looking ahead, portions of the analysis can migrate to lightweight on-device models to further reduce reliance on server-side processing.

### v2

@@ 4.3. Server Architecture and Processing Pipeline
The computational demands of this server processing pipeline are substantial. For example, processing a single Reddit batch can require 50 to 250 sequential external LVLM API calls that must complete within seconds to maintain usability. In our deployment, the critical path for processing an initial 25-post Reddit batch completed in roughly 5–15 seconds, with subsequent batches typically processed before users scrolled to them. We achieved real-time performance through asynchronous processing with multi-batch responses, content-based caching, and predictive prefetching. We have described these optimizations and additional design decisions in Appendix˜A.
@@ 4.4. The Intervention Palette
The design of our system is based on a shift from content suppression to transformation. In taking this approach, we faced a design challenge: every intervention has to navigate a trade-off between three competing goals. The first is semantic fidelity, or how faithful the modified content is to the original’s intended meaning. The second is trigger fidelity, or how close the modified distressing element is to its original form. The final goal is perceptual smoothness, or how natural the result of the intervention looks to the user. We designed a palette of transformations, each reflecting a different approach to these trade-offs. Across this palette, interventions range from softening distressing elements to removing them, so people can still engage with the valuable parts of a content without being exposed to their triggers at full intensity. This approach echoes exposure therapy’s emphasis on working with weakened versions of stimuli.
@@
In contrast, semantic modifications prioritize perceptual smoothness. Inpainting(yu2018inpainting; li2024mat) as an example case, can seamlessly remove an object and reconstruct the background. This is a greater sacrifice of semantic fidelity than occlusion because it actively creates a new, plausible reality where the object never existed, rather than simply occluding part of the original. Another semantic intervention, visual euphemism, explores a different trade-off. It replaces a triggering object with a benign alternative, for instance, substituting spider with a leaf. While this also sacrifices semantic fidelity, the goal is to produce a visually coherent image that we hypothesize is less distressing to the user. To personalize this, users can add additional metadata to their filters or give examples to guide the visual euphemism choices, which may in turn alter the mood of the post.
Third, we also explore how altering an image’s rendering style could offer a middle ground. For example, reducing photorealism with artistic alteration, such as  Pointillism (clement1999neo) can lower the trigger fidelity of a sensitive region while preserving more structural information than occlusion. For these alterations, we borrow the aesthetic language of well-known artistic styles—Pointillism, Cubism, Impressionism, and Studio Ghibli style animation. We intended for each artistic style to not only reduce realism, but also act as an affective vessel that shapes how content is experienced. Impressionism, with its soft brushstrokes and diffused light, often evokes a sense of tranquility, distance, or dream-like calm. Studio Ghibli inspired style, with its rounded and whimsical rendering, lends even serious scenes a tone of cuteness and warmth. Cubism, which fragments forms in abstract planes, introduces a kind of conceptual distance, perhaps inviting interpretation over emotional reaction.
@@
In DIY-MOD, we implemented a representative set of interventions that explore the broader design space of content transformation. Figure˜2 illustrates several examples from our image palette’s three main categories. We list all implemented interventions in Table˜1 & 2 . The range of transformation options presents an important question: which intervention is most appropriate for a specific user and post? Next, we detail the framework we developed to address this question.
Category Intervention Description & Relation to Design Space
Warning
Overlay
Stylistic Alteration*
Can apply to both sensitive region only or whole image
Intervention Description & Relation to Design Space
Overlay
Warning
Hides an entire text block behind a personalized warning message.
@@ 4.5. Intervention Selection Framework
We developed a two-stage pipeline (Figure˜3) that evaluates interventions through the lens of individual user needs while maintaining real-time performance. Content that matches a user’s filters enters Stage 1 for pruning, and Stage 2 then generates and scores candidate transformations before selecting the best one.
@@ 4.5.1. Evaluation Criteria:
The three dimensions of semantic fidelity, trigger fidelity, and perceptual smoothness characterize the design space, but do not, on their own, determine which intervention is best suited for a given scenario. That choice requires translating these properties into outcomes relevant to a specific user’s safety and context. To guide this selection and evaluate outcomes, we operationalize a framework that decomposes intervention quality into four complementary dimensions, assessed through LLM-as-a-judge (zheng2023judging; liu2023gpteval):
  1. (1)
    Transformation Seamlessness: This dimension evaluates the perceptual smoothness of the output or the coherence of text. It maps directly to the perceptual smoothness axis of our design space, ensuring the intervention is not jarring.
  2. (2)
    Semantic Fidelity: Inherited directly from our design space, this assesses whether original facts and context are preserved without adding misinformation or hallucinations, a known risk in generative models (huang2025survey; gao2024aigcs).
  3. (3)
    Predicted Emotional Safety: For each user–content pair, this dimension acts as a personalized proxy for trigger fidelity. It evaluates how well a given transformation, in light of that user’s specific context (filter description, sensitivity level, chat history, and metadata preferences), is expected to reduce trigger exposure and meet the user’s safety needs.
  4. (4)
    Contextual Harm Risk: This dimension functions as a safeguard, assessing the degree to which a transformation might trivialize or distort themes that individuals or broader culture consider sensitive. For instance, applying Studio Ghibli style to war imagery or racial violence would would produce a high contextual-harm score. This risk score is subtracted from the overall intervention rating, so transformations that produce contextually inappropriate content are strongly penalized even if they score well on other dimensions.
@@ 4.5.2. Two-Stage Selection Pipeline
Stage 1: Predictive Pruning. The Pruner uses GPT-4o to perform multi-task inference(son2024multi) that first checks whether content matches any user filters. For matching content, it predicts which interventions would most likely succeed for that specific user. We provide the Pruner LVLM with detailed descriptions of each intervention including trade-offs, advantages, disadvantages, and ideal use cases. We also supply our four scoring dimensions and instruct the Pruner to evaluate how each intervention would score. This process approximates the full scoring process without generating actual transformations and produces a ranked list from which we select the top-$K$ candidates (typically $K=3$).
Stage 2: Generation and Scoring. DIY-MOD applies the top-$K$ candidates to the content, in parallel. Obfuscation techniques are applied locally using GroundingDino (liu2023grounding) and Python Pillow (clark2015pillow) library. For semantic modifications and stylistic alterations, we use Gemini’s asynchronous image generation model: gemini-2.0-flash-preview-image-generation (google_gemini2_0flash_2025). The Scorer uses GPT-4o-2024-08-06 to evaluate each candidate transformation against the evaluation dimensions (Section˜4.5.1). To mitigate potential biases inherent in LLM-as-a-judge systems where a model may favor its own outputs, we deliberately employ a different model for the scoring phase (GPT-4o) than for the generation phase (Gemini 2.0). We implement two guardrails in the scoring process. First, the Scorer receives both the original and transformed content, and is instructed to penalize misapplied interventions and failures. Second, we employ two-stage prompting (wang2024rl) that improve judge-VLM’s performance and results in more reliable and consistent evaluations. The Scorer first performs chain-of-thought analysis along our evaluation dimensions given the original content, transformed candidate, and user context. Based on this free-form response, it then produces a final numerical score. These scoring operations run in parallel across all $K$ candidates, and the highest-scoring intervention is selected.
@@ 5.1. Methods
Data Collection and Analysis: We transcribed all think-aloud data and interviews, then analyzed them using reflexive thematic analysis (braun2019reflecting) to identify patterns in user experiences and system interactions. During consent, we explicitly warned participants they might encounter content matching their stated sensitivities and emphasized they could pause, skip content, or withdraw at any point.
@@ 6.1. Method
As participants scrolled through the dual feed, they encountered two distinct, modified versions of each post with text and image transformation applied ( Figure˜4). For each pair, they were asked to select the version that better met their safety needs and explain their reasoning. The two modifications were chosen systematically: one was the intervention ranked highest by DIY-MOD’s selection framework, while the other was randomly selected from the framework’s other high-ranking alternatives (ranks 2-5). To mitigate positional bias, the on-screen placement (left or right) of these two versions was randomized for each post. While the primary task involved comparing the two modified versions, the original unmodified content was also available for context. Participants could request to view it or have it described verbally if needed. The study lasted approximately 30 minutes for each participant.
@@ 6.2.3. Quantitative Alignment with User Preferences
To complement our qualitative findings, we modeled the probability that participants would select the system’s highest-ranked intervention over a nearby alternative—the randomly selected intervention from ranks two through five. We use an intercept-only model to predict our dependent variable, System Choice Favored, which is a binary variable equal to 1 if the participant selected the system’s choice, and 0 otherwise. We use the glmer function from the lme4 package(lme4) in R. For image interventions, the estimated intercept was positive and statistically significant (log-odds = 0.53, $p=.033$), corresponding to a 63% probability that participants chose the system’s top candidate over the close contenders. For text interventions, the alignment was stronger (log-odds = $1.33$, p<.001{p}<.001), corresponding to a 79.1% probability.
@@ 7.1. Transformation Enables Access, Not Avoidance
Empowering users to take moderation decisions into their own hands might raise concerns about whether this agency can trap them in filter bubbles(pariser2011filter). DIY-MOD can support a different dynamic. By transforming how content appears rather than removing it entirely, the system enables users to engage with content they would otherwise avoid completely.
@@
At the same time, this approach surfaces deeper tensions: Who decides what users should or should not engage with, and why does that decision need to be made in a top-down, prescriptive manner? The First Amendment is often framed as a protection of speech. Legal scholars have similarly argued for the right to freedom of listening and the right against compelled listening (corbin2009first). While these protections do not extend to platforms as private entities, the principles that they represent remain relevant. Our studies show that content that appears benign to most can be experienced as harmful by others. This variability complicates any universal judgment about what content is “safe” or “appropriate”—at least outside the categories codified by law. And it raises a critical question: Who gets to decide which experiences of harm are valid and which users are entitled to protection?
@@ 7.2. The Case for Platform-Integrated Personalization
While DIY-MOD demonstrates personalized moderation is viable as middleware (fukuyama2021), its architecture requires intercepting, externally processing, and then scoring generated transformations. These steps introduce unavoidable latency and complicate implementation.
Platforms however, could avoid these bottlenecks entirely. With direct content access, computational resources, and existing user behavioral signals, they could implement personalized moderation far more efficiently. Platforms already use this infrastructure for ad targeting and safety interventions like suicide prevention (facebook_suicide_1; facebook_suicide_2). The same systems could also power consensual, user-directed personalized content transformation.
@@
Our approach gives rise to a natural question: Which kinds of content should be addressed through platform-led moderation, and which might be better suited to user-driven transformation? Platforms are widely expected to bear responsibility for addressing harms that have broad, collective consequences such as coordinated manipulation campaigns or abuse like hate speech or targeted harassment that fosters hostile environments. Yet even here, the boundaries are far from settled. The same activity can be cast as harassment or as activism, as disinformation or as political dissent. Platforms’ judgments about what to enforce are never neutral. They are shaped by political pressures, business incentives, and cultural norms (diaz2021double; jahanbakhsh2023thesis).
@@ 7.4. Application to Civic Discourse
This transformation-based approach may also hold promise for other types of content not brought up by our study participants. One can imagine its application to civic or ideological discourse: when valuable perspectives are obscured by hostile or inflammatory framing, repackaging that content (e.g., by softening antagonistic language, reducing confrontational cues, or highlighting shared values), could make it easier to engage with. It is also worth noting that there is no clear consensus that exposure to counter-attitudinal information is always beneficial. In fact, a body of research shows that such exposure, especially when unfiltered or confrontational, can exacerbate polarization and deepen animosity (tornberg2022digital_echo; bail2022breaking). This calls into question that more exposure is always better, and highlights the need for systems that support constructive engagement, rather than unmediated confrontation (bail2022breaking). Altering the tone or phrasing of civic or political content through transformations like ours may offer a path toward openness without overwhelm. It is also worth considering who is most likely to adopt such tools. Prior work suggests challenge-averse individuals may use personalization to reinforce existing views (munson2010presenting). However, those findings reflect systems that offer a binary choice (e.g., show/hide). By contrast, transformation reduces the affective load of counter-attitudinal content. It is plausible that this, unlike binary filtering, might enable engagement from users who would otherwise disengage completely. Investigating these adoption patterns remains a critical area for future research.
Prior work on bridging divides has largely focused on identifying and ranking content that resonates across ideological lines (ovadya2022bridging; stray2021designing). While valuable, such approaches depend on the availability of naturally bridgeable posts. But what if we could make more content bridgeable? This of course, raises important questions about authenticity and message integrity: How much transformation is too much? When does reframing a message become misrepresentation? Future work must address these ethical boundaries.
@@ 7.5. Therapeutic Grounding and Applications of Our Work
DIY-MOD’s design is inspired by exposure therapy: enabling safer engagement with a weakened or modified version of a stimulus, rather than avoidance (rothbaum2002exposure; abramowitz2019exposure). This principle is reflected in our palette of interventions. Our transformations are designed to function as weakened versions of the original content, and the ability to easily adjust sensitivity levels and filter durations echoes the concept of graduated exposure.
While DIY-MOD is not a clinical intervention in its current form, its combination of weakened content, adjustable sensitivities, and time-bounded filters hints at how similar mechanisms could eventually complement therapeutic practices. For example, these knobs could let users (and clinicians) gradually decrease intervention intensity over time, if they so choose, rather than assuming static sensitivities. This direction aligns with calls (wang2025mentally) for practitioner-assisted AI systems that support personalized, adaptive exposure in everyday contexts.
@@ 7.6.1. Architectural Safeguards:
Using AI to interpret personal sensitivities carries inherent risks, as system failures can expose users to content they explicitly sought to avoid. One way we mitigate these risks is by implementing a two-stage cascade architecture: a pruner and then a scorer VLM-judge evaluate content using a consistent rubric. Such multi-stage evaluation strengthens the reliability of LLM/VLM-judge based pipelines (badshah2024reference; gu2024survey). We manually audited sampled content and found that all sampled text edits and approximately 89% of sampled image edits were successfully transformed, with false obfuscations (edits that did not match the viewer’s filter description) in only about 7% of transformed images (Section˜A.5).
@@
Privacy considerations are fundamental when handling data tied to personal sensitivities. DIY-MOD adopts a “privacy-by-design” approach (cavoukian2009privacy) that prioritizes user control and data minimization. It provides full functionality without an account and offers user-controlled portability of filters through an anonymous export/import feature. By default, it uses local-first storage and practices data minimization, sending only public in-feed content for ephemeral analysis. Looking ahead, portions of the analysis can migrate to lightweight on-device models to further reduce reliance on server-side processing.